OpenStudy (calculusxy):

MEDAL for help in geometry! Prove the following. Given: \(\angle 3\) is supp. to \(\angle 1\); \(\angle 4\) is supp. to \(\angle 2\). Prove: \(\angle 3\) \(\cong\) \(\angle 4\)

OpenStudy (calculusxy):

@zepdrix @agent0smith

OpenStudy (calculusxy):



OpenStudy (mayankdevnani):

Suppose \[\large \bf \angle 3=x\] then \[\large \bf \angle1=180-x\] similarly, \[\large \bf \angle4=y\] then \[\large \bf \angle 2=180-y\] And we know that \(\large \bf \angle1~and~\angle2\) are vertically opposite angles. therefore, \[\large \bf \angle1=\angle2\] plug the values in it, we get \[\large \bf 180-x=180-y\] \[\large \bf \color{red}{x}=\color{red}{y}\] therefore, \[\large \bf \angle3~\cong~\angle 4\]

OpenStudy (mayankdevnani):

Hope it helps :D

OpenStudy (calculusxy):

I need to write a two-coloumn proof

OpenStudy (mayankdevnani):